OpenStudy (anonymous):

what is the geometric mean of 6 and 7

jimthompson5910 (jim_thompson5910):

Rule: \[\Large \text{Geometric mean of x and y} = \sqrt{x*y}\]

OpenStudy (mathstudent55):

The geometric mean of a and b is x, when they are in this proportion: \(\dfrac{a}{x} = \dfrac{x}{b} \) When you solve for x, you get: \(x = \sqrt{ab} \)